OpenStudy (anonymous):

y varies directly as the square of x. when x=4, y=8. find x when y= 32

OpenStudy (anonymous):

y=kx^2 k=y/x^2 =8/4 =1/2

OpenStudy (anonymous):

when y = 32 x^2 = y/k => x^2 =16 =>x = 4

OpenStudy (anonymous):

x=4